\begin{thebibliography}{99}
\bibitem{nbs}M. Abramowitz and I. A. Stegun,
{Handbook of Mathematical Functions, Washington, National Bureau of Standards
(1964).}
%\bibitem{addison}A. W. Addison,
%{A series representation for Euler's constant, Amer. Math. Monthly {\bf 74}, 
%823-824 (1967).}
\bibitem{andrews}G. E. Andrews, R. Askey, and R. Roy, 
{Special Functions, Cambridge University Press (1999).}
%could add Koekek's errata list for A^2R too ...
%\bibitem{apostol}T. M. Apostol,
%{Introduction to Analytic Number Theory, Springer Verlag, New York (1976);
%corrected fourth printing, 1995.}
%\bibitem{baez}L. B\'{a}ez-Duarte,
%{On Maslanka's representation for the Riemann zeta function, preprint (2003);
%A new necessary and sufficient condition for the Riemann hypothesis, preprint 
%(2003).}
\bibitem{bb07}D. H. Bailey et al.,
{Experimental Mathematics in Action, A. K. Peters, Wellesley, MA (2007).}
\bibitem{bbnotices}D. H. Bailey and J. M. Borwein,
{Experimental mathematics:  Examples, methods and implications, Notices Amer. Math. Soc.
{\bf 52}, 502-514 (2005).}
\bibitem{bbcomp}D. H. Bailey and J. M. Borwein,
{Computer-assisted discovery and proof, in:  Tapas in Experimental Mathematics,
Contemp. Math., T. Amdeberhan and V. Moll, eds., Amer. Math. Soc. (2008), pp.
21-52; preprint 
\newline{http://crd.lbl.gov/$\sim$dhbailey/dhbpapers/comp-disc-proof.pdf (2007).}}
%to appear some place?--send note to Dave Bailey about ......
%\bibitem{baileyetal}D. H. Bailey, J. M. Borwein, D. Broadhurst, and W. Zudilin, {Experimental %mathematics and mathematical physics, to appear in: More Tapas in Experimental Mathematics, volume %xx of Contemp. Math.;  http://www.carma.newcastle.edu.au/$\sim$jb616/bbbz09.pdf (2009).}
\bibitem{berndt}B. C. Berndt,
{On the Hurwitz zeta-function, Rocky Mtn. J. Math {\bf 2}, 151-157 (1972).}
\bibitem{bb98}J. M. Borwein and D. J. Broadhurst,
{Determination of rational Dedekind-zeta invariants of hyperbolic manifolds and
Feynman knots and links, arxiv:hep-th/9811173 (1998).}  %still unpublished?
%\bibitem{balazard}M. Balazard, E. Saias, and M. Yor,
%{Notes sur la fonction $\zeta$ de Riemann, 2, Advances in Mathematics 
%{\bf 143}, 284-287 (1999).}
%\bibitem{brun}V. Brun,
%{Deux transformations \'{e}l\'{e}mentaires de la fonction Zeta de Riemann, Rivista Ci. Lima {\bf %41}, 517-525 (1939).}
\bibitem{selberg67}S. Chowla and A. Selberg,
{On Epstein's Zeta-function, J. Reine Angew. Math. {\bf 227}, 86-110 (1967).}
%\bibitem{coffeyjnt}M. W. Coffey,
%{Addison-type series representation for the Stieltjes constants, J. Number Th. {\bf 130}, 2049-2064 %(2010); arXiv:0912.2391 (2009).}
%\bibitem{coffey04}M. W. Coffey,
%{Relations and positivity results for derivatives of the Riemann $\xi$
%function, J. Comput. Appl. Math., {\bf 166}, 525-534 (2004).}
%\bibitem{coffey03}M. W. Coffey,
%{Towards verification of the Riemann hypothesis:  Application of the
%Li criterion, preprint, 2003.}
%\bibitem{coffeygam1}M. W. Coffey,
%{New results on the Li constants, preprint, 2004.}
\bibitem{coffeyjmaa}M. W. Coffey,
{New results on the Stieltjes constants:  Asymptotic and exact evaluation, 
J. Math. Anal. Appl. {\bf 317}, 603-612 (2006); arXiv:math-ph/0506061.}
%\bibitem{coffeyrama}M. W. Coffey,
%{Functional equations of the Stieltjes constants, Ramanujan J. {\bf 39}, 577-601 (2016).}
%%\bibitem{coffey04}M. W. Coffey,
%%{The Li constants are asymptotically positive, and more, preprint, 2004.}
%\bibitem{coffeyhecke}M. W. Coffey,
%{An explicit formula and estimations for Hecke L functions, preprint, 2004.}
%\bibitem{cofunpub}M. W. Coffey,
%{unpublished (2004).}
%\bibitem{cofpla}M. W. Coffey,
%{A set of identities for a theta function at unit argument, Phys. Lett. A,
%{\bf 300}, 367-369 (2002).}
\bibitem{coffeyrecip}M. W. Coffey,
{Parameterized summation relations for the Stieltjes constants, arXiv:1002.4684v2 (2010).}
\bibitem{coffeystdiffs}M. W. Coffey,
{On representations and differences of Stieltjes coefficients, and other relations, Rocky Mtn. J. Math. {\bf 41}, 1815-1846 (2011); arXiv/math-ph/0809.3277v2 (2008).}
\bibitem{coffey2009}M. W. Coffey,  
{Series representations for the Stieltjes constants, Rocky Mtn. J. Math. {\bf 44}, 443-477 (2014), arXiv:0905.1111 (2009).}
\bibitem{coffeygammaratl}M. W. Coffey,     %UPDATE w/ publ info:
{Functional equations for the Stieltjes constants, Ramanujan Journal, {\bf 39}, 577-601, (2016), arXiv:1402.3746 (2014).}
\bibitem{coffeyjmp}M. W. Coffey,
{Evaluation of a $\ln \tan$ integral arising in quantum field theory, J. Math. Phys. 
{\bf 49}, 093508-1-15 (2008).}
\bibitem{coffeyjmp2}M. W. Coffey,
{On a three-dimensional symmetric Ising tetrahedron, and contributions
to the theory of the dilogarithm and Clausen functions, J. Math. Phys. {\bf 49}, 
043510-1-32 (2008).}
\bibitem{coffeyanly}M. W. Coffey,
{The Stieltjes constants, their relation to the $\eta_j$ coefficients, and
representation of the Hurwitz zeta function, Analysis {\bf 30}, 383-409 (2010); arxiv/math.ph/0706343 (2007).}
%\bibitem{coffeystv2}M. W. Coffey,
%{The Stieltjes constants, their relation to the $\eta_j$ coefficients,
%and representation of the Hurwitz zeta function, arXiv/math-ph/:0706.0343v2 %(2007).}
%\bibitem{coffeyprsa}M. W. Coffey,
%{New results on power series expansions of the Riemann xi function and the
%Li/Keiper constants, Proc. Royal Soc. A {\bf 464}, 711-731 (2008).}
%\bibitem{davenport}H. Davenport,
%{Multiplicative Number Theory, Springer Verlag (2000).}
\bibitem{debruijn}N. G. de Bruijn,
{Integralen voor de $\zeta$-function van Riemann, Mathematica (Zutphen) {\bf B5}, 170-180 (1937).}
\bibitem{edwards}H. M. Edwards,
{Riemann's Zeta Function, Academic Press, New York (1974).}
\bibitem{gordon}L. Gordon,
{A stochastic approach to the Gamma function, Amer. Math. Monthly {\bf 101}, 858-865 (1994).}
%\bibitem{gould}H. W. Gould,
%{Euler's formula for nth differences of powers, Amer. Math. Monthly
%{\bf 85}, 450-467 (1978).}
\bibitem{grad}I. S. Gradshteyn and I. M. Ryzhik,
{Table of Integrals, Series, and Products, Academic Press, New York (1980).}
%\bibitem{hansen}E. R. Hansen and M. L. Patrick,
%{Some relations and values for the generalized Riemann zeta function,
%Math. Comp. {\bf 16}, 265-274 (1962).}
%\bibitem{hua}L. K. Hua,
%{Introductiion to Number Theory, Springer Verlag (1982).}
%\bibitem{israilov}M. I. Israilov,
%{On the Laurent Decomposition of Riemann's zeta function,
%Dokl. Akad. Nauk SSSR (Russian) {\bf 12}, 9 (1979); M. I. Israilov,
%Trudy Mat. Inst. Steklova {\bf 158}, 98-104 (1981).}
\bibitem{ihara}Y. Ihara,
{On the Euler-Kronecker constants of global fields and primes with small norms,
in:  Algebraic Geometry and Number Theory, ed.\ V.\ Ginzburg, Birkh\"{a}user (2006).}
\bibitem{ireland}K. Ireland and M. Rosen,
{A Classical Introduction to Modern Number Theory, 2nd ed., Springer (1990).}
\bibitem{ivic}A. Ivi\'{c}, 
{The Riemann Zeta-Function, Wiley, New York (1985).}
%\bibitem{ivic93}A. Ivi\'{c},
%{The Laurent coefficients of certain Dirichlet series,
%Publs. Inst. Math. {\bf 53}, 23-36 (1993).  In Eq. (1.7), $(\log x)^x$ should
%be replaced with $(\log x)^k$ and on p. 25 "comparing with (1.15)" should be 
%replaced with "comparing with (1.5)".}
%\bibitem{jensen}J. L. W. V. Jensen,
%{Sur la fonction $\zeta(s)$ de Riemann, C. R. Acad. Sci. Paris {\bf 104}, 1156-1157
%(1887).}
%\bibitem{kanemitsu}S. Kanemitsu, H. Kumagai, and M. Yoshimoto,
%{Sums involving the Hurwitz zeta function, Ramanujan J. {\bf 5}, 5-19 (2001).}
%\bibitem{karatsuba}A. A. Karatsuba and S. M. Voronin,
%{The Riemann Zeta-Function, Walter de Gruyter, New York (1992).}
%\bibitem{keiper}J. B. Keiper,
%{Power series expansions of Riemann's $\xi$ function, Math. Comp. {\bf 58}, 
%765-773 (1992).}
\bibitem{knessl2}C. Knessl and M. W. Coffey,
{An asymptotic form for the Stieltjes constants $\gamma_k(a)$ and for a sum $S_\gamma(n)$ appearing under the Li criterion Math. Comp. {\bf 80}, 2197-2217 (2010).}
%\bibitem{knopp}K. Knopp,
%{Theory and Application of Infinite Series, Hafner Publishing Company, New
%York (1951).}
%\bibitem{lachaud}G. Lachaud,
%{Spectral analysis and the Riemann hypothesis, J. Comp. Appl. Math., {\bf 160},
%175-190 (2003).  On p. 176 of this paper, $\xi(s)$ is missing the factors
%$s(s-1)/2$ and on p. 178 the definition of ${\cal L}$ is misstated.}
%\bibitem{garunk}A. Laurin\v{c}ikas and R. Garunk\v{s}tis,
%{The Lerch Zeta-function, Kluwer Academic Publishers Dordrecht (2002).}
%\bibitem{hlmont}H. L. Montgomery and A. M. Odlyzko,
%{Large deviations of sums of independent random variables, Acta Arith.
%{\bf XLIX}, 427-434 (1988).}
\bibitem{mathar}R. J. Mathar,
{Table of Dirichlet $L$-series and prime zeta modulo functions for small moduli,
arXiv:1008.2547 (2010).}
%\bibitem{mati}Yu. V. Matiyasevich,
%{A relationship between certain sums over trivial and nontrivial zeros of the
%Riemann zeta function, Mat. Zametki {\bf 45}, 65-70 (1989).}
%\bibitem{neukirch}J. Neukirch,
%{Algebraic Number Theory, Springer Verlag (1999).}
%\bibitem{odlyzko}A. M. Odlyzko,
%{http://www.dtc.umn.edu/\~odlyzko/} %put something else in for this reference?
%\bibitem{amo1989}A. M. Odlyzko,
%{Bounds for discriminants and related estimates for class numbers, regulators 
%and zeros of zeta functions:  A survey of recent results, Sem. Th. No., 
%Bordeaux, {\bf g}, 1-15 (1989).}
%\bibitem{omar}S. Omar,
%{Localization of the first zero of the Dedekind zeta function, Math. Comp.
%{\bf 70}, 1607-1616 (2001).}
%\bibitem{olver}F. W. J. Olver,
%{Asymptotics and Special Functions, Academic Press (1974).}
\bibitem{paris}R. B. Paris,
{An asymptotic expansion for the Stieltjes constants, arXiv:1508.03946v2 (2015).} % now published?
\bibitem{riemann}B. Riemann,
{\"{U}ber die Anzahl der Primzahlen unter einer gegebenen Gr\"{o}sse, 
Monats. Preuss. Akad. Wiss., 671 (1859-1860).}
%\bibitem{rsmith}R. Smith,
%{private communication (2008).}
\bibitem{sierpref}W. Sierpinski,
{On the summation of the series $\sum_{n>a}^{n \leq b}\tau(n)f(n)$, where $\tau(n)$ denotes the
number of decompositions of $n$ into a sum of two integer squares (in Polish), Prace
Matematyczno-Fizyczne {\bf 18}, 1-59 (1908); French transl. in Oeuvres Choisies, t. 1, Editions
Scientifiques de Pologne, 109-154 (1974).}
\bibitem{stieltjes}T. J. Stieltjes,
{Correspondance d'Hermite et de Stieltjes, Volumes 1 and 2, Gauthier-Villars,
Paris (1905).}
%\bibitem{szego}G. Szeg\"{o},
%{Orthogonal Polynomials, Vol. 23 of AMS Colloquium Publications, AMS (1975).}
%look up this reference:
%\bibitem{temme}N. Temme,
%{Asymptotic estimates for Laguerre polynomials, J. Appl. Math. Phys. {\bf 41},
%114-126 (1990).}  %also to look up:
\bibitem{titch}E. C. Titchmarsh,
{The Theory of the Riemann Zeta-Function, 2nd ed., Oxford University
Press, Oxford (1986).}
%\bibitem{watson}E. T. Whittaker and G. N. Watson,
%{A course of modern analysis, 4th ed., Cambridge Univ. Press, Cambridge (1962).}
%\bibitem{weil}A. Weil,
%{Basic Number Theory, Springer Verlag (1967).}
%\bibitem{weil}A. Weil,
%{Oeuvres Scientifiques--Collected Papers, Vol. II, pp. 48-61, Springer
%(1980).}
%\bibitem{eric}E. W. Weisstein, 
%{"Stieltjes Constants." From MathWorld--A Wolfram Web Resource. %http://mathworld.wolfram.com/StieltjesConstants.html} 
\bibitem{wilton2}J. R. Wilton, 
{A note on the coefficients in the expansion of $\zeta(s,x)$ in powers of 
$s-1$, Quart. J. Pure Appl. Math. {\bf 50}, 329-332 (1927).}
%\bibitem{wilton}J. R. Wilton,
%{A proof of Burnside's formula for $\log \Gamma(x+1)$ and certain allied properties 
%of Riemann's $\zeta$-function, Messenger Math. {\bf 52}, 90-93 (1922-1923).}
\bibitem{zagiergangl}D. Zagier and H. Gangl, 
{Classical and elliptic polylogarithms and special values of $L$-series, The arithmetic and geometry of algebraic cycles, Banff, AB, 1998, NATO Science Series C, Mathematical and Physical Sciences, {\bf 548}, 561--615, Kluwer Academic Publishers, Dordrecht (2000).}
\bibitem{zucker76}I. J. Zucker and M. M. Robertson,
{Some properties of Dirichlet $L$-series, J. Phys. A {\bf 9}, 1207-1214 (1976).}
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphicx}
\usepackage{graphics}
\usepackage{amsthm}
\usepackage{amssymb, amsmath}
\title{Summatory relations and prime products for the Stieltjes constants, and other related results} 
\author{Mark W. Coffey\\
Department of Physics\\
Colorado School of Mines\\
Golden, CO  80401\\
(Received $\mbox{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~2016}$)}
\date{July 31, 2016}
\begin{document}
\maketitle
\baselineskip=25 pt
\begin{abstract}

The Stieltjes constants $\gamma_k(a)$ appear in the regular part of the Laurent expansion for the Hurwitz zeta function $\zeta(s,a)$.  We present summatory results for these constants $\gamma_k(a)$ in terms of fundamental mathematical constants such as the Catalan constant, and further
relate them to products of rational functions of prime numbers.  
We provide examples of infinite series of differences of Stieltjes constants evaluating as
volumes in hyperbolic $3$-space.  We present a new series representation for the difference
of the first Stieltjes constant at rational arguments.  
We obtain expressions for $\zeta(1/2)L_{-p}(1/2)$, where for primes $p>7$, $L_{-p}(s)$ are certain
$L$-series, and remarkably tight bounds for the value $\zeta(1/2)$, $\zeta(s)=\zeta(s,1)$ being
the Riemann zeta function.




\end{abstract}
 
\medskip
\baselineskip=15pt
\centerline{\bf Key words and phrases}
\medskip 

\noindent

Dirichlet $L$ function, Hurwitz zeta function, Stieltjes constants, series representation, Riemann zeta function, Laurent expansion, digamma function, summatory relation, prime product

\vfill
\centerline{\bf 2010 AMS codes} 
11M06, 11Y60, 11M35

\baselineskip=25pt
\pagebreak
\medskip
\centerline{\bf Introduction}
\medskip

Let $\zeta(s,a)$ be the Hurwitz zeta function, and $\zeta(s,1)=\zeta(s)$ be the Riemann
zeta function \cite{edwards,ivic,riemann,titch} .  In the complex plane of $s$, each of these functions has a simple pole of residue $1$ at $s=1$.  For Re $s>1$ and Re $a>0$ we have
$$\zeta(s,a)=\sum_{n=0}^\infty {1 \over {(n+a)^s}}, \eqno(1.1)$$
and by analytic continuation $\zeta(s,a)$ extends to a meromorphic function through out
the whole complex plane. Accordingly, there is the Laurent expansion in terms of the
Stieltjes constants $\gamma_k(a)$,
\cite{berndt,coffeyjmaa,coffeystdiffs,coffey2009,stieltjes,wilton2}
$$\zeta(s,a)={1 \over {s-1}}+\sum_{n=0}^\infty {{(-1)^n} \over {n!}}\gamma_n(a)(s-1)^n,
~~~~~~s \neq 1.  \eqno(1.2)$$
We let $\Gamma(s)$ be the Gamma function, $\psi(s)=\Gamma'(s)/\Gamma(s)$ be the digamma function (e.g., \cite{nbs,andrews,grad}) with the Euler constant $\gamma=\gamma_0(1)=-\psi(1)$, and recall that $\gamma_0(a)=-\psi(a)$. In the following, $\psi^{(k)}$ are the polygamma functions.
We note that for integers $n>1$, $\zeta(n,a)$ reduces to values $\psi^{(n-1)}$.

The sequence $\{\gamma_k(a)\}_{k=0}^\infty$ exhibits complicated changes in
sign with $k$.  For instance, for both even and odd index, there are infinitely
many positive and negative values.  Furthermore, there is sign variation with the
parameter $a$.  These features, as well as the exponential growth in magnitude
in $k$, are now well captured in an asymptotic expression (\cite{knessl2}, Section 2, \cite{paris}).
In fact, though initially derived for large values of $k$, this expression is
useful for computational approximation even for small values of $k$.

Reciprocity and other relations for the Stieltjes constants, Bernoulli polynomials, and
functions $A_k(q)$ are given in \cite{coffeyrecip}.  
The works \cite{coffeygammaratl} and \cite{coffey2009} present explicit expressions for
the initial values of the Stieltjes constants at rational argument, and various series
representations, respectively.  


This paper is largely illustrative, being a step in a program to explore summatory
relations for the Stieltjes constants in terms of fundamental mathematical constants such 
as the Catalan constant
$$G \equiv \sum_{n=0}^\infty {{(-1)^n} \over {(2n+1)^2}}.  \eqno(1.3)$$  
We also relate sums of differences of Stieltjes constants to other mathematical constants 
and products of prime numbers.  
\footnote{For such prime product results, especially see Proposition 7}.
For the latter purpose, $p$ is reserved for the product index
over primes.  Two of many examples to be presented are then
$$\sum_{k=0}^\infty {1 \over {k!}}\left[\gamma_k\left({1 \over 3}\right)-\gamma_k\left({2 \over 3}\right)\right]={1 \over 3}, \eqno(1.4a)$$
and
$$G={\pi^2 \over 8}\prod_{\overset{p \equiv 3}{mod ~4}} {{p^2-1} \over {p^2+1}}={1 \over {16}}
\sum_{k=0}^\infty {1 \over {k!}}\left[\gamma_k\left({1 \over 4}\right)-\gamma_k\left({3 \over 4}\right)\right]. \eqno(1.4b)$$
While, for instance, the first equality in (1.4b) is known, it seems unlikely that the
connection with the Stieljtes constants in the second equality has been established before.


A major tool for the first sections of this paper are quadratic Dirichlet-$L$ series and
their properties.  Therefore, some further notation and definitions are introduced.  Let $D$
be a fundamental discriminant and $(D/n)$ the Kronecker-Jacobi-Legendre symbol.
This symbol, a completely multiplicative function on the positive integers, is a real primitive Dirichlet character with modulus $|D|$.  Then the Dirichlet
$L$ series associated to $(D/n)$ is defined for Re $s>1$ as
$$L_D(s)=\sum_{n=1}^\infty \left({D \over n}\right)n^{-s}. \eqno(1.5)$$
If $D=1$, then $L_1(s)=\zeta(s)$.  For all other values of $D$, $L_D(s)$ can be made into an
entire function by using the value $L_D(1) \neq 0$.  
With suitable factors of $D^{s/2}$, $\pi^{-s/2}$, and $\Gamma[(s+\epsilon)/2]$ ($\epsilon=0,1$),
$L_D(s)$ may be completed to $L^*_D(s)$ with the compact functional equation 
$L^*_D(s)=L_D^*(1-s)$.  A useful compilation of special values of $L_D(s)$ is contained in
\cite{mathar}.

Importantly for what follows, we describe how Dirichlet $L$-functions $L_{\pm k}(s)$ (e.g., \cite{ireland}, Ch. 16), are expressible as linear combinations of Hurwitz zeta functions.  
We let $\chi_k$ be a real Dirichlet character modulo $k$, where the corresponding $L$ 
function is written with subscript $\pm k$ according to $\chi_k(k-1)=\pm 1$.  We have
$$L_{\pm k}(s)=\sum_{n=1}^\infty {{\chi_k(n)} \over n^s}={1 \over k^s}\sum_{m=1}^k 
\chi_k(m) \zeta\left(s,{m \over k}\right), ~~~~\mbox{Re} ~s >1.  \eqno(1.6)$$
This equation holds for at least Re $s>1$.  If $\chi_k$ is a nonprincipal character, 
as we typically assume in the following, then convergence obtains for Re $s>0$.

These $L$ functions, extendable to the whole complex plane, satisfy the functional
equations \cite{zucker76}
$$L_{-k}(s)={1 \over \pi}(2\pi)^s k^{-s+1/2}\cos\left({{s \pi} \over 2}\right)\Gamma
(1-s)L_{-k}(1-s), \eqno(1.7)$$
and
$$L_{+k}(s)={1 \over \pi}(2\pi)^s k^{-s+1/2}\sin\left({{s \pi} \over 2}\right)\Gamma
(1-s)L_{+k}(1-s). \eqno(1.8)$$
Owing to the relation
$$\Gamma(1-s)\Gamma(s)={\pi \over {\sin \pi s}}, \eqno(1.9)$$
these functional equations may also be written in the form
$$L_{-k}(1-s)=2(2\pi)^{-s} k^{s-1/2}\sin\left({{\pi s} \over 2}\right) \Gamma(s)
L_{-k}(s), \eqno(1.10)$$
and
$$L_{+k}(1-s)=2(2\pi)^{-s} k^{s-1/2}\cos\left({{\pi s} \over 2}\right) \Gamma(s)
L_{+k}(s). \eqno(1.11)$$


%

\medskip
\centerline{\bf Statement of results}
\medskip

The first result may be placed in the context of special function theory, while the
subsequent ones Propositions 2--8 and 10 provide summatory results for differences of the Stieltjes constants.  Proposition 9 concerns representation of Sierpinski's constant $S$,
and the last Proposition 11 bounds $\zeta(s)$ in the critical strip for real values of $s$.
\newline{\bf Proposition 1} (a) 
$$L_{\pm k}(1)={1 \over k}\sum_{m=1}^k\chi_k(m)\gamma_0\left({m \over k}\right),$$
and (b) for all $D \neq 1$, the values
$$L_D(1)= \begin{cases}
{\pi \over {3\sqrt{3}}} ~\mbox{if}~ D=-3 \\
{\pi \over 4} ~\mbox{if}~ D=-4 \\
{{\pi h(D)} \over \sqrt{-D}} ~\mbox{if}~ D<-4 \\
{{2 h(D)\ln \varepsilon} \over \sqrt{D}} ~\mbox{if}~ D>1 \\
\end{cases}, $$ 
may be written in terms of $\gamma_0=-\psi$. 
Here $h(D)$ is the ideal class number of the quadratic field $\mathbb{Q}(\sqrt{D})$ and
$\varepsilon$ is the fundamental unit of the integer subring $\mathbb{Z}+((D+\sqrt{D})/2)\mathbb{Z}$
.

{\bf Proposition 2} (a) 
$$L_{\pm k}(2)={1 \over k^2}\sum_{n=0}^\infty {{(-1)^n} \over {n!}} \sum_{m=1}^k\chi_k(m)\gamma_n\left({m \over k}\right),$$
and (b) for all $D$, the values $L_D(2)$ 
may similarly be written in terms of sums of differences $\gamma_k(a)$. 
This includes the cases $L_1(2)=\zeta(2)=\pi^2/6$, $L_{-4}(2)=G$, Catalan's constant,
$$L_5(2)={{4\pi^2} \over {25\sqrt{5}}}, ~~~~L_8(2)={\pi^2 \over {8\sqrt{2}}}, ~~~~
L_{12}(2)={\pi^2 \over {6\sqrt{3}}},$$
and
$$L_{-7}(2)=I_{-7} \equiv {{24} \over {7 \sqrt{7}}}\int_{\pi/3}^{\pi/2} \ln \left|{{\tan t+\sqrt{7}} \over {\tan t-\sqrt{7}}}\right |dt. \eqno(2.1)$$

The value (2.1) and related integrals arise in hyperbolic geometry, knot theory, and quantum field
theory, and (2.1) has received considerable attention in the last several years \cite{bb07,bbnotices,bbcomp,bb98}.
References \cite{coffeyjmp,coffeyjmp2} provide alternative evaluations of the integral (2.1)
in terms of the Clausen function Cl$_2(\theta)$, 
$$\mbox{Cl}_2(\theta)\equiv -\int_0^\theta\ln\left|2 \sin {t \over 2}\right|dt
=\int_0^1 \tan^{-1}\left({{x \sin \theta} \over {1-x\cos \theta}}\right)
{{dx} \over x}$$
$$=-\sin \theta \int_0^1 {{\ln x} \over {x^2-2x\cos \theta +1}}dx
=\sum_{n=1}^\infty {{\sin(n \theta)} \over n^2}.$$
The evaluations of Proposition 2 provide connections of infinite sums of differences of Stieltjes constants with values Cl$_2(\theta)$.

The next result, among other relations, connects a sum of differences of Stieltjes constants with
a sum involving the sum of divisors function $\sigma_{-3}(n)=\sum_{d|n}d^{-3}$.  In stating the
result, we introduce the Bloch-Wigner dilogarithm
$$D(z)=\mbox{Im}[\mbox{Li}_2(z)+\ln |z|\ln(1-z)]= \mbox{Im}[\mbox{Li}_2(z)]+\mbox{arg}(1-z) \ln |z|,$$
where, as usual for $|z|\leq 1$, the dilogarithm function is given by
$\mbox{Li}_2(z)=\sum_{n=1}^\infty {z^n \over n^2}$, which analytically continues to $\mathbb{C}
\backslash (1,\infty)$. 
Then $D(z)$ is real analytic in $\mathbb{C}\backslash \{0,1\}$.
\newline{\bf Proposition 3} Let $z_1=(1+i\sqrt{23})/2$, $z_2=2+i\sqrt{23}$, $z_3=(3+i\sqrt{23})/2$, $z_4=
(5+i\sqrt{23})/2$, and $z_5=3+i\sqrt{23}$.  Then
$${3 \over 2}\zeta(4)+{{40\pi} \over 23^{3/2}}\left[{{\zeta(3)} \over 2}+\sum_{n=1}^\infty
\sigma_{-3}(n)(1+\pi n\sqrt{23})e^{-\pi n\sqrt{3}}\right]$$
$$={{4\pi^2/3} \over {23^{3/2}}}[21D(z_1)+7D(z_2)+D(z_3)-3D(z_4)+D(z_5)]$$
$$=23^{-2}\sum_{n=0}^\infty {{(-1)^n} \over {n!}}\sum_{j=1}^{22}\chi_{-23}(j)\gamma_n\left({j \over
{23}}\right)$$
$$\equiv 23^{-2}\sum_{n=0}^\infty {{(-1)^n} \over {n!}}\left[\gamma_n\left({1 \over {23}}\right)+
\gamma_n\left({2 \over {23}}\right)+\gamma_n\left({3 \over {23}}\right)+
\gamma_n\left({4 \over {23}}\right)-\gamma_n\left({5 \over {23}}\right)+
\gamma_n\left({6 \over {23}}\right) \right.$$
$$-\gamma_n\left({7 \over {23}}\right)+
\gamma_n\left({8 \over {23}}\right)+\gamma_n\left({9 \over {23}}\right)-
\gamma_n\left({{10} \over {23}}\right)+\gamma_n\left({{11} \over {23}}\right)+
\gamma_n\left({{12} \over {23}}\right)-\gamma_n\left({{13} \over {23}}\right)+\gamma_n\left({{14} \over {23}}\right)$$
$$\left. -\gamma_n\left({{15} \over {23}}\right)+\gamma_n\left({{16} \over {23}}\right)
-\gamma_n\left({{17} \over {23}}\right)+\gamma_n\left({{18} \over {23}}\right)-
\gamma_n\left({{19} \over {23}}\right)+\gamma_n\left({{20} \over {23}}\right)-\gamma_n\left({{21} \over {23}}\right)-\gamma_n\left({{22} \over {23}}\right)\right].$$

{\bf Proposition 4} (a) 
$$L_{\pm k}(3)={1 \over k^3}\sum_{n=0}^\infty {{(-1)^n} \over {n!}}2^n \sum_{m=1}^k\chi_k(m)\gamma_n \left({m \over k}\right),$$
and (b) for all $D$, the values $L_D(2)$ 
may similarly be written in terms of sums of differences $\gamma_k(a)$. 
In particular, for $D<0$, closed-form expressions for $L_D(3)$ are known.  

After the proof of Proposition 4 we illustrate the use of an integral representation for
$\gamma_k(a)$ in order to obtain such sums for rational values of $a$.

{\bf Proposition 5} 
$$4^{-4} \sum_{n=0}^\infty{{(-1)^n} \over {n!}}3^n\left[\gamma_n\left({1 \over 4}\right)-
\gamma_n\left({3 \over 4}\right)\right]={1 \over {768}}\psi'''\left({1 \over 4}\right)-{\pi^4 \over {96}}.$$

The next sort of result follows from consideration of the Euler-Kronecker constant $\gamma_D$
of the quadratic field $\mathbb{Q}(\sqrt{D})$.  It provides a new summation representation of
the difference of the first Stieltjes constant at rational arguments.
\newline{\bf Proposition 6} (a)
$$\gamma_1\left({3 \over 4}\right)-\gamma_1\left({1 \over 4}\right)={\pi^2 \over 3}+\pi \gamma
+4\pi \sum_{\ell=1}^\infty {1 \over \ell}{1 \over {(e^{2\pi \ell}-1)}},$$
and (b)
$$\gamma_1\left({2 \over 3}\right)-\gamma_1\left({1 \over 3}\right)={\pi \over \sqrt{3}}\left[
{\pi \over {2\sqrt{3}}}+\gamma -4\sum_{\ell=1}^\infty {{(-1)^\ell} \over \ell}{1 \over {[(-1)^\ell-e^{\sqrt{3}\pi \ell}]}}\right].$$

As we mention below, such differences of $\gamma_1(a)$ at rational argument are related to 
values of $\ln \Gamma(a)$ and hence $\ln \Gamma(1-a)$.  For further details \cite{coffeygammaratl}
may be consulted.

{\bf Proposition 7} (a)
$$\prod_{\overset{p \equiv 1}{mod ~3}} {{p^2+1} \over {p^2-1}}={3 \over {2\pi^2}}
\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\left[\gamma_k\left({1 \over 3}\right)-\gamma_k\left({2 \over 3}\right)\right],$$
$${4 \over {27}}\prod_{\overset{p \equiv 2}{mod ~3}} {{p^2-1} \over {p^2+1}}={1 \over 9}
\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\left[\gamma_k\left({1 \over 3}\right)-\gamma_k\left({2 \over 3}\right)\right],$$
(b) (1.4b) holds, as well as
$$\prod_{\overset{p \equiv 1}{mod ~4}} {{p^2+1} \over {p^2-1}}={{12} \over \pi^2}G={4 \over {3\pi^2}} \sum_{k=0}^\infty {1 \over {k!}}\left[\gamma_k\left({1 \over 4}\right)-\gamma_k\left(
{3 \over 4}\right)\right],$$
(c)
$${1 \over \sqrt{5}}=\prod_{\overset{p \equiv 2}{mod ~5}} {{p^2-1} \over {p^2+1}}
\prod_{\overset{p \equiv 3}{mod ~5}} {{p^2-1} \over {p^2+1}}$$
$$=\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\left[\gamma_k\left({1 \over 5}\right)-
\gamma_k\left({2 \over 5}\right)-\gamma_k\left({3 \over 5}\right)+
\gamma_k\left({4 \over 5}\right)\right],$$
$${{124} \over {125}}\zeta(3)= \prod_{\overset{p \equiv 2}{mod ~5}} {{p^3-1} \over {p^3+1}}
\prod_{\overset{p \equiv 3}{mod ~5}} {{p^3-1} \over {p^3+1}}$$
$$={1 \over {125}}\sum_{k=0}^\infty {{(-2)^k} \over {k!}}\left[\gamma_k\left({1 \over 5}\right)-
\gamma_k\left({2 \over 5}\right)-\gamma_k\left({3 \over 5}\right)+
\gamma_k\left({4 \over 5}\right)\right]$$
$$={1 \over {125}}\left[\psi''\left({2 \over 5}\right)-
\psi''\left({1 \over 5}\right)+\psi''\left({3 \over 5}\right)-
\psi''\left({4 \over 5}\right)\right],$$
(d) 
$${{3\pi^3} \over {64\sqrt{2}}}= {7 \over 8}\zeta(3)\prod_{\overset{p \equiv 5}{mod ~8}} {{p^3-1} \over {p^3+1}} \prod_{\overset{p \equiv 7}{mod ~8}} {{p^3-1} \over {p^3+1}}$$
$$={1 \over {8^3}}\sum_{k=0}^\infty {{(-2)^k} \over {k!}}\left[\gamma_k\left({1 \over 8}\right)+
\gamma_k\left({3 \over 8}\right)-\gamma_k\left({5 \over 8}\right)-
\gamma_k\left({7 \over 8}\right)\right],$$
$${{\pi^2} \over {\sqrt{2}}}= 6\zeta(2)\prod_{\overset{p \equiv 3}{mod ~8}} {{p^2-1} \over {p^2+1}} \prod_{\overset{p \equiv 5}{mod ~8}} {{p^2-1} \over {p^2+1}}$$
$$={1 \over {8}}\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\left[\gamma_k\left({1 \over 8}\right)-
\gamma_k\left({3 \over 8}\right)-\gamma_k\left({5 \over 8}\right)+
\gamma_k\left({7 \over 8}\right)\right],$$
and (e)
$$\sum_{k=0}^\infty {{(-1)^k} \over {k!}}\left[\gamma_k\left({1 \over 8}\right)+
\gamma_k\left({3 \over 8}\right)-\gamma_k\left({5 \over 8}\right)-
\gamma_k\left({7 \over 8}\right)\right]
={{45} \over {32}}{{\zeta(4)} \over {\sqrt{2}G}}\prod_{\overset{p \equiv 7}{mod ~8}} \left({{p^2-1} \over {p^2+1}}\right)^2,$$
$$\sum_{k=0}^\infty {{(-2)^k} \over {k!}}\left[\gamma_k\left({1 \over 8}\right)-
\gamma_k\left({3 \over 8}\right)-\gamma_k\left({5 \over 8}\right)+
\gamma_k\left({7 \over 8}\right)\right]
={{3\pi^6} \over {1792\sqrt{2}}}{1 \over {\zeta(3)}}\prod_{\overset{p \equiv 7}{mod ~8}} \left({{p^3-1} \over {p^3+1}}\right)^2.$$


Let $E_j$ denote the Euler numbers, 
for which the initial values are $E_0=1$, $E_2=-1$, $E_4=5$, $E_6=-61$, $E_8=1385$,
$E_{10}=-50521$, and $E_{2n+1}=0$ for $n \geq 0$.

The next result complements Proposition 2 in the case that $D=-4$.
\newline{\bf Proposition 8}  
Let $k \geq 0$.  Then
$${{(-1)^k} \over {2(2k)!}}E_{2k}\left({\pi \over 2}\right)^{2k+1}=\sum_{n=0}^\infty {{(-1)^n} 
\over {n!}}\left[\gamma_n\left({1 \over 4}\right)-\gamma_n\left({3 \over 4}\right)\right](2k)^n.$$

{\bf Proposition 9}  
Sierpinski's constant \cite{sierpref}
$S=\gamma_{-4}\simeq 0.8228252496$ has the representation $S=2\gamma+{4 \over \pi}J_2$, wherein
$$J_2 \equiv \int_0^1 {{\ln(-\ln x)} \over {1+x^2}}dx$$
has the novel representation for $0<b<\pi/2$
$$J_2={1 \over 2}\sum_{k=1}^\infty {{E_{2k}b^{2k+1}} \over {(2k+1)!}}\left[\ln b -{b \over {(2k+1)}}\right]+{b \over 2}(\ln b-1)+{1 \over 2}\int_b^\infty {{\ln u} \over {\cosh u}}du.$$
In particular,
$$J_2=-{1 \over 2}\left[1+\sum_{k=1}^\infty {{E_{2k}} \over {(2k+1)!}}{1 \over {(2k+1)}}\right]
+{1 \over 2}\int_1^\infty {{\ln u} \over {\cosh u}}du.$$

{\bf Proposition 10}.  Let $\sigma_1(n)$ be the sum of divisors function and $K_0(z)$ the
zeroth order modified Bessel function of the second kind.  Then
(a)
$$\zeta\left({1 \over 2}\right)L_{-11}\left({1 \over 2}\right)=\gamma+\ln\left({\sqrt{11} \over
{8\pi}}\right)+4\sum_{n=1}^\infty (-1)^n \sigma_1(n)K_0(\sqrt{11}n\pi)$$
$$=\zeta\left({1 \over 2}\right){1 \over \sqrt{11}}\sum_{n=0}^\infty {1 \over 2^n}{1 \over {n!}}
\sum_{k=1}^{10}\left({k \over {11}}\right)\gamma_n\left({k \over {11}}\right),$$
and, more generally, (b) for $p >7$ a prime and class number $h(-p)=1$,
\footnote{so that $p>7$ is given by $11$, $19$, $43$, and $67$.}
$$\zeta\left({1 \over 2}\right)L_{-p}\left({1 \over 2}\right)=\gamma+\ln\left({\sqrt{p} \over
{8\pi}}\right)+4\sum_{n=1}^\infty (-1)^n \sigma_1(n)K_0(\sqrt{p}n\pi)$$
$$=\zeta\left({1 \over 2}\right){1 \over \sqrt{p}}\sum_{n=0}^\infty {1 \over 2^n}{1 \over {n!}}
\sum_{k=1}^{p-1}\left({k \over p}\right)\gamma_n\left({k \over p}\right).$$

{\bf Proposition 11}.  Let $0<s<1$.  Then $\zeta(s)<0$.  In particular, $\zeta(1/2)<0$.
In fact we also show that 
$$-{3 \over 2}+{1 \over {15 \sqrt{5}}}<\zeta\left({1 \over 2}\right)<-{{35} \over {24}}.$$

\medskip
\centerline{\bf Proof of Propositions}

{\it Proposition 1}.  (a) This follows from (1.6) and then (1.2),
$$L_{\pm k}(1)=\sum_{n-1}^\infty {{\chi_k(n)} \over n}={1 \over k} \sum_{m=1}^k \chi_k(m)
\zeta\left(1,{m \over k}\right)$$
$$={1 \over k}\sum_{m=1}^k\chi_k(m)\gamma_0\left({m \over k}\right),$$
(b) follows since $\sum_{n=1}^k \chi_k(n)=0$ for nonprincipal characters $\chi_k$.

{\it Proposition 2}. This again follows from (1.6) and the use of (1.2).

{\it Remark}.  From the relation $L_{-7}(2)=I_{-7}$ and \cite{coffeyjmp} we have for instance
$${1 \over \sqrt{7}}\sum_{n=0}^\infty{{(-1)^n} \over {n!}}\left[\gamma_n\left({1 \over 7}\right)+
\gamma_n\left({2 \over 7}\right)-\gamma_n\left({3 \over 7}\right)+
\gamma_n\left({4 \over 7}\right)-\gamma_n\left({5 \over 7}\right)-
\gamma_n\left({6 \over 7}\right)\right]$$
$$=4\left[3\mbox{Cl}_2(\theta_7)-3\mbox{Cl}_2(2\theta_7)+\mbox{Cl}_2(3\theta_7)\right].$$

{\it Proposition 3}.  The Dedekind zeta function of $\mathbb{Q}(\sqrt{-23})$ is given by
$$\zeta_{\mathbb{Q}(\sqrt{-23})}(s)=\zeta_0(s)+2\zeta_1(s)$$
$$=\sum_{m,n}'\left[{1 \over 2}{1 \over {(m^2+mn+6n^2)^s}}+{1 \over {(2m^2+mn+3n^2)^s}}\right],$$
with the prime on the sum indicating the exclusion of the term $m=n=0$.
We apply (1.2), (1.6), and the values \cite{zagiergangl}
$$\zeta_{\mathbb{Q}(\sqrt{-23})}(2)={{4\pi^2/3} \over {23^{3/2}}}[21D(z_1)+7D(z_2)+D(z_3)-3D(z_4)+D(z_5)]$$
$$=\zeta_0(2)+2\zeta_1(2),$$
where the partial zeta function values are given by
$$\zeta_0(2)=\zeta(4)+{{8\pi} \over {23^{3/2}}}+\left[{{\zeta(3)} \over 2}+\sum_{n=1}^\infty
\sigma_{-3}(n)(1+\pi n\sqrt{23})e^{-\pi n\sqrt{3}}\right],$$
and
$$\zeta_1(2)={{\zeta(4)} \over 4}+{{16\pi} \over {23^{3/2}}}+\left[{{\zeta(3)} \over 2}+\sum_{n=1}^\infty \sigma_{-3}(n)(1+\pi n\sqrt{23})e^{-\pi n\sqrt{3}}\right].$$

{\it Remark}.  Proposition 3 gives an example of the general relation for volumes in hyperbolic
$3$-space $\mathbb{H}^3$
$$\zeta_F(2)={{4\pi^2} \over D^{3/2}}\mbox{Vol}\left(\mathbb{H}^3/SL_2({\cal{O}}_F)\right),$$
where $F$ is an imaginary quadratic field.

{\it Proposition 4}. The results follow from (1.6) and the use of (1.2).

{\it Remark}.  In particular, we have corresponding summations of differences of Stieltjes
constants for values including
$$L_{-3}(3)={{4 \pi^3} \over {81\sqrt{3}}}, ~~~~L_{-4}(3)={\pi^3 \over {32}},$$
$$L_{-7}(3)={{32 \pi^3} \over {343\sqrt{7}}}, ~~~~L_{-8}(3)={{3\pi^3} \over {64\sqrt{2}}}.$$
For example,
$${{4 \pi^3} \over {81}}={1 \over 3^{5/2}}\sum_{n=0}^\infty{{(-1)^n} \over {n!}}2^n\left[\gamma_n\left({1 \over 3}\right)-\gamma_n\left({2 \over 3}\right)\right],$$
and
$${{32 \pi^3} \over {343}}={1 \over 7^{5/2}}\sum_{n=0}^\infty{{(-1)^n} \over {n!}}2^n\left[\gamma_n\left({1 \over 7}\right)+
\gamma_n\left({2 \over 7}\right)-\gamma_n\left({3 \over 7}\right)+
\gamma_n\left({4 \over 7}\right)-\gamma_n\left({5 \over 7}\right)-
\gamma_n\left({6 \over 7}\right)\right].$$

Therefore we have identities for the summation of differences of Stieltjes constants as
given by the following example.
\newline{\bf Corollary 1}.
$$\pi^3={{81} \over {4\cdot 3^{5/2}}}\sum_{n=0}^\infty{{(-1)^n} \over {n!}}2^n\left[\gamma_n\left({1 \over 3}\right)-\gamma_n\left({2 \over 3}\right)\right]$$
$$={{343} \over {32 \cdot 7^{5/2}}}\sum_{n=0}^\infty{{(-1)^n} \over {n!}}2^n\left[\gamma_n\left({1 \over 7}\right)+
\gamma_n\left({2 \over 7}\right)-\gamma_n\left({3 \over 7}\right)+
\gamma_n\left({4 \over 7}\right)-\gamma_n\left({5 \over 7}\right)-
\gamma_n\left({6 \over 7}\right)\right].$$

Another systematic way to produce summatory relations for differences of the Stieltjes
constants is to use the author's integral representation \cite{coffeyanly} 
for Re $a>0$,
$$\gamma_k(a)={1 \over {2a}}\ln^k a-{{\ln^{k+1} a} \over {k+1}}+{2 \over a}\mbox{Re}~
\int_0^\infty {{(y/a-i)\ln^k(a-iy)} \over {(1+y^2/a^2)(e^{2\pi y}-1)}}dy.$$
We then determine that, for instance,
$$\sum_{k=0}^\infty{{(-2)^k} \over {k!}}\left[\gamma_k\left({1 \over 3}\right)-\gamma_k\left({2 \over 3}\right)\right]={{243} \over {16}}$$
$$+54 \mbox{Re}\int_0^\infty\left[{1 \over {(3y+2i)^3}}-{1 \over {(3y+i)^3}}\right]{{dy} \over
{(e^{2\pi y}-1)}}dy.$$
For $q$ real, we observe that
$$\mbox{Re}~{1 \over {(3y+qi)^3}}=9y{{(3y^2-q^2)} \over {(9y^2+q^2)^3}}.$$
Then
$$\sum_{k=0}^\infty{{(-2)^n} \over {k!}}\left[\gamma_k\left({1 \over 3}\right)-\gamma_k\left({2 \over 3}\right)\right]={{243} \over {16}}
+456\int_0^\infty\left[{{3y^2-4} \over {(3y^2+4)^3}}-{{(3y^2-1)} \over {(3y^2+1)^3}}\right]{{ydy} \over {(e^{2\pi y}-1)}}.$$
By taking two derivatives of an integral representation for the digamma function, 
$$\psi(s)=\ln s-{1 \over {2s}}-2\int_0^\infty {{tdt} \over {(t^2+s^2)(e^{2\pi t}-1)}},$$
we obtain for the tetragamma function
$$\psi''(s)=-{1 \over s^2}-{1 \over s^3}+8\int_0^\infty {{(t^2-s^2)} \over {(t^2+s^2)^3}}{{tdt}
\over {(e^{2\pi t}-1)}}.$$
With the difference $\psi''(2/3)-\psi''(1/3)=8\pi^3/3^{3/2}$, we again obtain the result stated in the remark above.

{\it Proposition 5}. We use (1.2), (1.6), and the value
$$L_{-4}(4)=4^{-4}[\zeta(4,1/4)-\zeta(4,3/4)]={1 \over {1536}}\left[\psi'''\left({1 \over 4}
\right)-\psi'''\left({3 \over 4}\right)\right]$$
$$={1 \over 768}\psi'''\left({1 \over 4}\right)-{\pi^4 \over {96}}.$$
The latter relation follows from the functional equation 
$$\psi'''(1-z)-\psi'''(z)=-2\pi^4(2+\cos 2\pi z)\csc^4 \pi z.$$

{\it Proposition 6}. The Euler-Kronecker constant of $\mathbb{Q}(\sqrt{D})$ is defined as
$$\gamma_D=\gamma+{{L_D'(1)} \over {L_D(1)}}.$$
(a)  For $D=-4$ and $L_{-4}(1)=\pi/4$,
$$\gamma_{-4}=\gamma+{4 \over \pi}L_{-4}'(1)  
=\ln\left[2\pi e^{2 \gamma}{{\Gamma^2(3/4)} \over {\Gamma^2(1/4)}}\right],$$
Sierpinski's constant \cite{sierpref}.
With $L_{-4}(s)=4^{-s}[\zeta(s,1/4)-\zeta(s,3/4)]$,
$$L_{-4}'(s)=-(\ln 4)L_{-4}(s)+4^{-s}[\zeta'(s,1/4)-\zeta'(s,3/4)].$$
From (1.2),
$$\left[\zeta'(s,a)+{1 \over {(s-1)^2}}\right]_{s=1}=-\gamma_1(a),$$
giving
$$L_{-4}'(1)=-(\ln 4){\pi \over 4}+4^{-1}\left[\gamma_1\left({3 \over 4}\right)-
\gamma_1\left({1 \over 4}\right)\right],$$
and
$$\gamma_{-4}=\gamma-\ln 4+{1 \over \pi}\left[\gamma_1\left({3 \over 4}\right)-
\gamma_1\left({1 \over 4}\right)\right].$$
Now by the Kronecker limit formula \cite{ihara} we also have
$$\gamma_{-4}={\pi \over 3}-\ln 4+2 \gamma-4\sum_{k=1}^\infty \ln(1-e^{-2\pi k}).$$
By equating the two expressions for $\gamma_{-4}$ we obtain
$${1 \over \pi}\left[\gamma_1\left({3 \over 4}\right)-\gamma_1\left({1 \over 4}\right)\right]
={\pi \over 3}+\gamma-4\sum_{k=1}^\infty \ln(1-e^{-2\pi k}).$$
If we expand the logarithm, and interchange sums applying geometric series, we obtain
the stated result.

(b) For $D=-3$ we have
$$\gamma_{-3}=\gamma+{{L_{-3}'(1)} \over {L_{-3}(1)}}
=\gamma+{{3\sqrt{3}} \over \pi}L_{-3}'(1)  
=\ln\left[2\pi e^{2 \gamma}{{\Gamma^3(2/3)} \over {\Gamma^3(1/3)}}\right].$$
Since
$$L_{-3}'(s)=-(\ln 3)L_{-3}(s)+3^{-s}[\zeta'(s,1/3)-\zeta'(s,2/3)],$$
$$\gamma_{-3}=\gamma-\ln 3+{\sqrt{3} \over \pi}\left[\gamma_1\left({2 \over 3}\right)-
\gamma_1\left({1 \over 3}\right)\right].$$
Also
$$\gamma_{-3}={\pi \over {2\sqrt{3}}}-\ln 3+2\gamma -4\sum_{k=1}^\infty \ln |1-e^{-2\pi i\omega_{-3}k}|,$$
where $\omega_{-3}=(1+i\sqrt{3})/2$.
Again by expanding in the logarithmic sum and rearranging,
$$\sum_{k=1}^\infty \ln |1-e^{-2\pi i\omega_{-3}k}|=-\sum_{k=1}^\infty \sum_{\ell=1}^\infty
{{(-1)^{\ell k}} \over \ell} e^{-\sqrt{3}\pi k \ell}
=\sum_{\ell=1}^\infty {{(-1)^\ell} \over \ell}{1 \over {[(-1)^\ell-e^{\sqrt{3}\pi \ell}]}},$$
we obtain the stated result.

{\it Proposition 7}. 
(a) is demonstrated via (1.2) and the Euler product of $L_{-3}(s)=L(s,\chi_{-3})$, so that
$$\prod_{\overset{p \equiv 1}{mod ~3}} {{p^2+1} \over {p^2-1}}={{27L_{-3}(2)} \over {2\pi^2}},$$
and
$$\prod_{\overset{p \equiv 2}{mod ~3}} {{p^2+1} \over {p^2-1}}={{4\pi^2} \over {27L_{-3}(2)}}.$$

(b) is proved similarly using $L_{-4}(s)$ and (1.2), and in this case we exhibit more of the
details of the Euler products for this $L$ series and $\zeta(s)$.  The resulting prime products
are
$$\prod_{\overset{p \equiv 1}{mod ~4}} {{p^2+1} \over {p^2-1}}={{12G} \over \pi^2},$$
and
$$\prod_{\overset{p \equiv 3}{mod ~4}} {{p^2+1} \over {p^2-1}}={\pi^2 \over {8G}}.$$

We have
$$L_{-4}(s)=\prod_p (1-\chi_{-4}(p)p^{-s})^{-1}=\prod_{\overset{p \equiv 1}{mod 4}} (1-p^{-s})^{-1}
\prod_{\overset{p \equiv 3}{mod ~4}} (1+p^{-s})^{-1},$$
while
$$\zeta(s)=(1-2^{-s})^{-1}\prod_{\overset{p \equiv 1}{mod ~4}} (1-p^{-s})^{-1}
\prod_{\overset{p \equiv 3}{mod ~4}} (1-p^{-s})^{-1}.$$
Then
$$L_{-4}(s)=(1-2^{-s})^{-1}\zeta(s)\prod_{\overset{p \equiv 3}{mod ~4}} {{(1-p^{-s})} \over 
{(1+p^{-s})}}=(1-2^{-s})^{-1}\zeta(s)\prod_{\overset{p \equiv 3}{mod ~4}} {{(p^s-1)} \over 
{(p^s-1)}}.$$
Then $L_{-4}(2)=G$ and (1.2) are used.  

(c) uses (1.2), $L_5(s)=5^{-s}[\zeta(s,1/5)-\zeta(s,2/5)-\zeta(s,3/5)+\zeta(s,4/5)]$, 
its Euler product representation, 
$$L_5(s)=\prod_{\overset{p \equiv 1}{mod ~5}} (1-p^{-s})^{-1}
\prod_{\overset{p \equiv 2}{mod ~5}} (1+p^{-s})^{-1} \prod_{\overset{p \equiv 3}{mod 5}} (1+p^{-s})^{-1} \prod_{\overset{p \equiv 4}{mod ~5}} (1-p^{-s})^{-1},$$
and the relation of the latter to that of $\zeta(s)$, so that
$$L_5(s)={{\zeta(s)} \over {(1-5^{-s})}}\prod_{\overset{p \equiv 2}{mod ~5}} {{(1-p^{-s})} \over 
{(1+p^{-s})}}\prod_{\overset{p \equiv 3}{mod ~5}} {{(1-p^{-s})} \over {(1+p^{-s})}}.$$
Evaluations at $s=2$ and $s=3$ are then performed.

(d) uses (1.2), $L_{-8}(s)=8^{-s}[\zeta(s,1/8)+\zeta(s,3/8)-\zeta(s,5/8)-\zeta(s,7/8)]$, 
and $L_{+8}(s)=8^{-s}[\zeta(s,1/8)-\zeta(s,3/8)-\zeta(s,5/8)+\zeta(s,7/8)]$, 
their Euler product representations in relation to that of $\zeta(s)$, and respective evaluations 
at $s=3$ and $s=2$.

(e) evaluates $L_{-8}(2)$ and $L_{+8}(3)$ using (1.2) and the Euler products of $L_{\pm 8}(s)$
in terms of those of $L_{-4}(s)$, $\zeta(s)$, and $\zeta(2s)$.

{\it Proposition 8}.  This is an application of (1.2), (1.6), and the evaluation for $k\geq 0$
$$L_{-4}(2k+1)={{(-1)^k} \over {2(2k)!}}E_{2k}\left({\pi \over 2}\right)^{2k+1}.$$

{\it Proposition 9}.  Based upon the integral $J_2$ evaluated in the Discussion section of
\cite{coffeygammaratl}, we again find
$$S=\gamma-\ln 4+{1 \over \pi}\left[\gamma_1\left({3 \over 4}\right)-\gamma_1\left({1 \over 4}\right)\right].$$
Then we have the various integral representations
$$S=2\gamma+{4 \over \pi}\int_0^\infty {{e^{-x}\ln x} \over {1+e^{-2x}}}dx
=2\gamma+{2 \over \pi}\int_0^\infty {{\ln x ~dx} \over {\cosh x}}=2\gamma+{4 \over \pi}J_2.$$
By applying a generating function of the Euler numbers,
$$\mbox{sech} ~x=1+\sum_{k=1}^\infty {{E_{2k}} \over {(2k)!}}x^{2k}, ~~~~|x| < {\pi \over 2},$$
we obtain the stated form of $J_2$.

{\it Proposition 10}.  We outline the steps for the more general setting of part (b).
Under the hypotheses stated, according to \cite{selberg67} (p.\ 95)
$$\zeta\left({1 \over 2}\right)L_{-p}\left({1 \over 2}\right)=\gamma+\ln\left({\sqrt{p} \over
{8\pi}}\right)+2\sum_{n=1}^\infty (-1)^n \sigma_1(n)\int_0^\infty e^{-\pi n \sqrt{p}(y+1/y)/2} {{dy} \over y}$$
$$=\gamma+\ln\left({\sqrt{p} \over {8\pi}}\right)+4\sum_{n=1}^\infty (-1)^n \sigma_1(n)K_0(\sqrt{p}n \pi).$$
Then (1.2) and (1.6) are used.

{\it Remarks}.  Such values $L_p(1/2)$ are of particular interest since finding an instance
of $L_p(1/2)<0$ would be a violation of the generalized Riemann hypothesis. 

Of course from (1.2) we have
$$\zeta\left({1 \over 2}\right)=-2+\sum_{n=0}^\infty {\gamma_n \over {2^nn!}}$$
$$=-2+{1 \over \pi}\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^{1/2}}
\simeq -1.46035 <0,$$
where $\gamma_n(1)=\gamma_n$.  Among many other occurrences of the value $\zeta(1/2)$, it
enters the Madelung constant $M_2=4(\sqrt{2}-1)\zeta(1/2)L_{-4}(1/2)$.  Hence we have
$$M_2=2(\sqrt{2}-1)\zeta\left({1 \over 2}\right)\sum_{n=0}^\infty {1 \over {2^n n!}}\left[
\gamma_n\left({1 \over 4}\right)-\gamma_n\left({3 \over 4}\right)\right].$$

{\it Proposition 11}.  For $0<\mbox{Re}~s<1$ \cite{debruijn}, 
$$\zeta(s)={1 \over {s-1}}+{{\sin\pi s} \over \pi}\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^s}.$$
Now for $x>0$,
$${1 \over {2x}}<\ln x-\psi(x)<{1 \over x}.$$
Letting $0<s<1$,
$${1 \over 2}\int_0^\infty {{dt} \over {(1+t)t^s}}<\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^s}<\int_0^\infty {{dt} \over {(1+t)t^s}}.$$
Then using a special case of the Beta function integral for Re$(a+b)>1$ and Re $b<1$,
$$\int_0^\infty {{dt}\over {(1+t)^a t^b}}=B(1-b,a+b-1)={{\Gamma(1-b)\Gamma(a+b-1)} \over
{\Gamma(a)}},$$
$${1 \over 2}<{{\sin\pi s} \over \pi}\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^s}<1.$$
Then it follows that $\zeta(s)<0$.

We may considerably refine the last inequality by using the inequalities from \cite{gordon}
(Theorem 5)
$${1 \over {2x}}+{1 \over {12(x+1/4)^2}}<\ln x-\psi(x)<{1 \over {2x}}+{1 \over {12x^2}}.$$
Restricting for illustration to $s=1/2$,
$${1 \over 2}\int_0^\infty {1 \over t^s}\left[{1 \over {t+1}}+{1 \over {6(t+5/4)^2}}\right]dt
<\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^{1/2}}<\int_0^\infty {1 \over t^{1/2}}
\left[{1 \over {t+1}}+{1 \over {6(t+1)^2}}\right]dt.$$
Thus we obtain
$${1 \over 2}\left(1+{2 \over {15\sqrt{5}}}\right)<{1 \over \pi}\int_0^\infty [\ln(1+t)-\psi(1+t)]{{dt} \over t^{1/2}} <{{13} \over {24}}.$$
Hence 
$$-{3 \over 2}+{1 \over {15 \sqrt{5}}}<\zeta\left({1 \over 2}\right)<-{{35} \over {24}}$$
follows.

{\it Remarks}.  Similarly we obtain 
$$-{5 \over 6}+{1 \over {\sqrt{2}\cdot 30\cdot 5^{1/4}}}<\zeta\left({1 \over 4}\right)=-{4 \over 3} +\sum_{n=0}^\infty \left({3 \over 4}\right)^n {\gamma_n \over {n!}} < -{{13} \over {16}},$$
which relates to $\zeta(3/4)$ via
$$\zeta\left({3 \over 4}\right)={{\sqrt{2+\sqrt{2}}\Gamma(1/4)} \over {(2\pi)^{1/4}}}\zeta\left(
{1 \over 4}\right).$$

From the value of $\zeta'(1/2)$ we obtain the identity
$${1 \over \pi}\int_0^\infty [\ln(1+t)-\psi(1+t)]{{\ln t} \over t^{1/2}}dt=\sum_{n=0}^\infty
{\gamma_{n+1} \over {2^n n!}}=-2\mbox{Re}~i\int_0^\infty {{\ln(1-iy)} \over {(1-iy)^{1/2}
(e^{2\pi y}-1)}}dy.$$ 





\pagebreak

\end{document}